OpenStudy (anonymous):

A pendulum has a period of.80s on Earth. What is its period on Mars, where the acceleration of gravity is about 0.37 that on Earth? How would I go by answering this problem? Thank you for any help!

OpenStudy (anonymous):

T on earth = .80 s g on earth = 9.8m/s^2 T on Mars = ? g on Mars = .37 * 9.8m/s^2 = 3.626m/s^2 T = 2 pi * sqrt L/g .80s = 6.28 * (sqrt L / sqrt 9.8m/^2) (.80s * 3.13) / 6.28 = sqrt L = .3987 L = .16m T = 2pi * ( sqrt .16m / 3.626m/s^2) T = 6.28 * (.4/ 1.904) = 6.28 * 0.21 T = 1.32s

OpenStudy (anonymous):

Thank you so much! I appreciate it :)